OpenStudy (anonymous):

Please help me out. :) Solve: lim tan 2x / tan x x approaches 0

OpenStudy (anonymous):

The answer given by the calculator is 2

OpenStudy (anonymous):

how to solve?

OpenStudy (anonymous):

\[\frac{ \tan2x }{ tanx }=\frac{ \frac{ 2tanx }{ 1=\tan^{2}x } }{ tanx }=\frac{ 2tanx }{ tanx(1-\tan^{2}x }=\frac{ 2 }{ 1-\tan^{2}x }\] \[\lim_{x \rightarrow 0}\frac{ 2 }{ 1-\tan^{2}x }=\frac{ 2 }{ 1 }=2\]

OpenStudy (anonymous):

\[\tan2x=\frac{ 2tanx }{ 1-\tan^{2}x }\]

OpenStudy (anonymous):

wow man smart :P

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

Wow! Thank you. :) ^_^